# On Picard’s Theorem Via Nevanlinna Theory

### v5

@@ 5. Calculus Lemma and Logarithmic Derivative Lemma
Let be a non-compact complete Kähler manifold of complex dimension with non-negative Ricci curvature. Assume that is non-parabolic.
@@ 5.1. Calculus Lemma
We need the following Borel’s lemma (see, e.g., [24]):
Lemma 5.1 (Borel Lemma).
@@
The conclusion is true clearly for Next, we assume that Since is a non-decreasing function, then there exists a number such that The non-decreasing property of implies that the limit
@@
Now, we give the Calculus Lemma:
@@
Invoking Theorem 4.1, we obtain
@@
holds for all outside Therefore, we have the theorem proved by setting
As a result, we conclude that
@@ 5.2. Logarithmic Derivative Lemma
in a local holomorphic coordinate where is the inverse of Again, define the Nevanlinna’s characteristic function of by
@@
Next, we establish the following Logarithmic Derivative Lemma:
@@
Using Lemma 5.5 and disturbing we have for any there exists a subset of finite Lebesgue measure such that
@@ 6. Second Main Theorem and Defect Relation
Let be a non-compact complete Kähler manifold of complex dimension with non-negative Ricci curvature. Assume that is non-parabolic.
Set
@@
Let be a divisor, where are prime divisors. The reduced form of is defined as
@@ 6.1. Second Main Theorem
Let be a reduced divisor of simple normal crossing type. We write which is an irreducible decomposition of Equip every holomorphic line bundle with a Hermitian metric which induces a Hermitian metric on Since then we can assume that As a result, gives a volume form on Pick such that and Moreover, we define a singular volume form
on Let
be a differentiably non-degenerate meromorphic mapping. Set
@@
in the sense of currents, where is the reduced form of Thus, it yields that
@@
Let be a non-compact complete Kähler manifold with non-negative Ricci curvature. Let be a complex projective manifold of complex dimension not greater than that of Let be a reduced divisor of simple normal crossing type, where is a positive line bundle over Let be a differentiably non-degenerate meromorphic mapping. If is non-parabolic, then for any there exists a subset of finite Lebesgue measure such that
@@
The argument of proof is standard (see [14]), but for the completeness of this paper, we still give a proof. Since has only simple normal crossings, then there exist a finite open covering of and finitely many rational functions on such that are holomorphic on satisfying that
In addition, we can require that for On write
@@
Fix a one can take a local holomorphic coordinate near and a local holomorphic coordinate near such that
Put
@@
Again, set for Then, at :
That is, at :
Combining the above, we are led to
on Note that is bounded on whence it yields from that
By Jensen-Dynkin formula
Combining (6) with (7) and using Theorem 5.6, we obtain
@@
holds for all outside Thus, we conclude that
for all outside Combining this with (6.1), we prove the theorem. ∎
Let be a compact Riemann surface of genus Let be the holomorphic line bundle over defined by a point Note that is independent of the choice of It yields from that
@@
Let be distinct points in Let be a nonconstant meromorphic mapping. Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside
Let be the hyperplane line bundle over and whose Chern form is given by It yields from that
Corollary 6.3.
Let be hyperplanes in general position in Let be a differentiably non-degenerate meromorphic mapping. Then for any there exists a subset of finite Lebesgue measure such that
@@ 6.2. Defect Relation
For any divisor the defect and simple defect of with respect to are defined respectively by
Using the First Main Theorem, we see that
For two holomorphic line bundles over we define (see [7])
@@
Endowing with Kähler metric Let be the Hilbert-Schmidt norm of differential with respect to metrics Let denote the scalar curvature of Take a local holomorphic coordinate the Ricci curvature tensor of can be written as the form
where
Using the definition of scalar curvature, we have
In further, it yields that
@@
Let be a meromorphic mapping. If there exists a small number such that satisfies the growth-curvature condition
@@
Assume on the contrary that is differentiably non-degenerate. Then, we have as The growth-curvature condition leads to
Thus, we obtain due to Corollary 6.4, but which contradicts with ∎
Treat a compact Riemann surface of genus we have:
@@
For a meromorphic mapping
we have
with respect to metrics
@@
We obtain a Picard’s theorem:
@@
6.3. Picard’s Theorem
In the following, we investigate a more exact form of the Picard’s theorem given in Corollary 6.8. On we introduce the spherical distance
@@ 6.3. Picard’s Theorem
We use the Ahlfors-Shimizu’s characteristic function of :
Moreover, we define
It is not hard to show the First Main Theorem:
@@
Let be a non-compact complete Kähler manifold with non-negative Ricci curvature. Let be distinct values in Let be a nonconstant meromorphic function on If is non-parabolic, then for any there exists a subset of finite Lebesgue measure such that
@@
By
in the sense of distributions, we deduce that
Combining this with Corollary 6.3, it gives immediately
We re-define the simple defect of with respect to by
@@
Consequently, we derive a more exact Picard’s theorem:

### v6

@@ 5. Calculus Lemma and Logarithmic Derivative Lemma
Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature, of complex dimension
@@ 5.1. Calculus Lemma
Lemma 5.1 (Borel’s Lemma).
@@
The conclusion is clearly true for Next, we assume that Since is a non-decreasing function, then there exists a number such that The non-decreasing property of implies that the limit
@@
We establish the Calculus Lemma as follows.
@@
Invoking Lemma 4.1, we have
@@
holds for all outside Therefore, we have the theorem proved by setting ∎
@@ 5.2. Logarithmic Derivative Lemma
in a holomorphic local coordinate where is the inverse of Again, define the Nevanlinna’s characteristic function of by
@@
Next, we establish the following Logarithmic Derivative Lemma.
@@
By Lemma 5.5, we have for any there exists a subset of finite Lebesgue measure such that
@@ 6. Second Main Theorem and Defect Relation
Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature, of complex dimension Let be complex projective manifold of complex dimension not greater than Over one can put a positive holomorphic line bundle Let be a reduced divisor of simple normal crossing type. Recall that
@@
Let be a divisor, where are prime divisors. The reduced form of is defined by
@@ 6.1. Second Main Theorem
Write as the irreducible decomposition of Equiping every holomorphic line bundle with a Hermitian metric such that it induces the Hermitian metric on Pick such that with On we define a singular volume form
where
Let be a differentiably non-degenerate meromorphic mapping. Set
@@
in the sense of currents. Thus, it yields that
@@
Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature. Let be a complex projective manifold of complex dimension not greater than that of Let be a reduced divisor of simple normal crossing type, where is a positive line bundle over Let be a differentiably non-degenerate meromorphic mapping. Then for any there exists a subset of finite Lebesgue measure such that
@@
Since is of simple normal crossing type, then we can provide a finite open covering of and finitely many rational functions on for every such that are holomorphic on with that
Moreover, we may require for On write
@@
Fix any we can choose a holomorphic local coordinate near and a holomorphic local coordinate near such that
Set
@@
Moreover, put for Then, at :
Whence, at :
Put together the above, we are led to
on Since is bounded on and
then we obtain
By Lemma 3.1
Combining (5) with (6) and using Theorem 5.6, we have
@@
holds for all outside Hence, we conclude that
for all outside By this with (6.1), we prove the theorem. ∎
@@
Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature. Let be a complex projective manifold of complex dimension not greater than that of Let be a reduced divisor of simple normal crossing type, where is a positive line bundle over Let be a differentiably non-degenerate meromorphic mapping. Then for any there exists a subset of finite Lebesgue measure such that
@@ 6.2. Defect Relation
The simple defect of with respect to is defined by
By the First Main Theorem, we have
By the use of Theorem it is immediate that
@@
Let
denote the Ricci curvature tensor of Since is Kählerian, then
Again, denote by the scalar curvature of Note that is the trace of thus we have
It yields that
@@
Let be a meromorphic mapping. If there exists a number such that satisfies the growth condition
@@
Assume on the contrary that is differentiably non-degenerate. Then, we obtain as The growth condition implies that
Thus, we deduce due to Corollary 6.3, but which contradicts with ∎
Consider a compact Riemann surface of genus Note that
for an arbitrary point where is the holomorphic line bundle over defined by Combining this with Theorem 6.1, we conclude that
@@
Let
be a meromorphic mapping. Then
Let be the hyperplane line bundle over Since
then Theorem 6.1 yields that
@@
We obtain a Picard’s theorem as follows.
@@
In the following, we shall derive a better Picard’s theorem compared with Corollary 6.7. On we introduce the spherical distance
@@
Adopt the Ahlfors-Shimizu’s characteristic function of as follows
Define
Then, we have the First Main Theorem
@@
Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature. Let be distinct values in Let be a nonconstant meromorphic function on Then for any there exists a subset of finite Lebesgue measure such that
@@
Define
@@
We derive a better Picard’s theorem as follows.
@@
Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature. Every meromorphic function on must be a constant if it omits distinct values, where denotes the maximal integer not greater than